OpenStudy (amistre64):

Let x,y \(\in~Z^+\) such that 11|(3x+7y). Which of these must also be divisible by 11? a. 4x+6y b. x+y+5 c. 9x+4y d. 4x - 9y e. x+y-1

OpenStudy (amistre64):

im confused as to why there would only be 1 correct choice

OpenStudy (amistre64):

prolly seeing it a little off, might need to run a euclidean/reverse on it to be sure

OpenStudy (amistre64):

7 = 3(2) + 1 0 0 1 0-2(1) = -2 ; x' = -2, y' = 1 11 * [3(-2)+7(1)=1] 3(-22)+7(11) = 11 x'' y'' x = x'' + 7n y = y'' - 3n

OpenStudy (primeralph):

not surre too.........